# Large cluster formation in a simple model for globular colloidal aggregates

### v2

@@ V Two-particle and intercluster structural correlations
Figure 10: Temperature and density dependence of the total structure factor S(Q) (main graph) and the intracluster structure factor (inset) for the system is stable cluster phases.
@@
If we now move from real to Fourier space, in Figure 9 we have the cluster-cluster structure factor, $S_{cl-cl}(Q)$. A first feature is to be noted, the height of the first peak for both densities is beyond Hansen-Verlet’sHansen and Verlet (1969) freezing rule $S(Q_{m})\gtrsim 2.86$ for which a transition towards a crystal is to be expected. As discussed above, our system presents features of an amorphous solid, but with a somewhat peculiar structure factor. We see that $S(Q)\approx 0$, $\forall\;Q\lesssim 0.2\sigma^{-1}$. In fact, one finds that the hyperuniformity index $H=S(Q_{max})/S(0<Q\lesssim 0.2\sigma^{-1})\approx 10^{3}\sim 10^{2}$ is relatively high. While its value is still away from the criterion for near hyperuniformity, $H\gtrsim 10^{4}$ proposed by Atkinson and coworkers Atkinson et al. (2016) in maximally random jammed systems, it represents a considerable attenuation of density fluctuations, and actually meets the criterion of effective hyperuniformity, $H\gtrsim 10^{2}$ proposed by Chen and et al.Chen et al. (2018). Lowering the temperature and choosing a material with the appropriate scattering form factor will improve the hyperuniform and stealthy character of the material for wavenumbers below $0.2\sigma^{-1}$. Tuning the size of the constituent particles one could increase or decrease the wavenumber range of stealthiness.
We note that the positions of the first maximum of the structure factors deviate from the value estimated from the correlation length, $Q_{m}=2\pi/\lambda_{0}$ (shown in the Figure by a vertical line), almost by 20%. Although the first peak of the intercluster pair distribution (Fig. 8) appears at $\approx\lambda_{0}$, the real effective cluster-cluster interaction is much softer than that depicted in the right graph of Figure 1. The particle density inside the clusters (cf. Figure 5) is far from the uniform density that was assumed in Eq. (4). This should lead to a softer outer range of the intercluster effective potential explaining the shift of $Q_{m}$ with respect to $2\pi/\lambda_{0}$. The fact that $Q_{m}(\rho\sigma^{3}=0.05)>Q_{m}(\rho\sigma^{3}=0.025)$ reflects the smaller size of the clusters for the lower density (see the corresponding radii of gyration in Figure 4) and matches the shift to lower $r$ of the first maximum of $g_{cl-cl}$ in Figure 8.
@@
If we now look at the total structural factor, $S(Q)$ depicted in Figure 10, its most noticeable feature is the huge magnitude of the prepeak, which is the signature of a large degree of clustering. Obviously the prepeak positions correspond to that of the first maxima in the cluster-cluster structure factors discussed above. Also here the hyperuniformity index meets the criterion of effective hyperuniformity.
@@
In the inset we can see the large wavenumber behavior of S(Q). The peak maxima $Q_{max}\sigma\approx 8.1$ reflects interparticle correlations at $0.8\sigma$, which is precisely the effective particle diameter, $\sigma_{eff}$. This is in accordance with the maxima of $g(r)$ in the inset of Figure 8. Interestingly, $H=S(Q_{max})/S(Q_{min})=0.2\times 10^{2}$ which is not very far from the criterion of effective hyperuniformity. A proper choice of the short range interparticle interactions might actually drive the system closer to having an additional wavenumber range where density fluctuations should be heavily attenuated.
@@ VI Conclusions
In summary, we have presented a study of the structural and dynamic properties of a simple model of a type-III SALR self-associating fluid, focusing on the globular cluster phase. Here we have studied both cluster size distributions, cluster density profiles, diffusion and frequency spectra as well as pair distributions (inter and intracluster) as well as structure factors. We have found that the cluster phase for moderate densities (still low packing fraction) as the temperature is lowered approaches a cluster glassy state. Interestingly, in these conditions the system shows a considerable attenuation of density fluctuations for a range of wavenumbers below $0.2\sigma^{-1}$, by which the material meets the criteria of effective hyperuniformity and stealthiness. Future work will focus on mixtures where intracluster phase separation can enable the tuning of anisotropic cluster-cluster interactions. At present, a two-dimensional counterpart of the system analyzed in this work is under study. An extension of the theoretical approach of Sweatman and coworkersSweatman et al. (2014) to include correlations at the HNC level so as to provide a more precise description of one and two-body cluster correlations is currently being developed.
Acknowledgements.
The authors acknowledge the support from the Agencia Estatal de Investigación and Fondo Europeo de Desarrollo Regional (FEDER) under grants No. PID2020-115722GB-C21 and PID2020-115722GB-C22. We also would like to acknowledge the Galicia Supercomputing Center (CESGA) for the access to their computer facilities.

### v4

@@ IV Single-cluster structure and dynamics
Figure 11: Density dependence of the total structure factor S(Q) (main graph) and the intracluster structure factor (upper right inset) for the system in stable cluster phases. In the lower left inset the low-Q behavior of the particle-particle structure factor is depicted.
We can now analyze two-particle correlations. To start with, in Figure 9 we present the cluster-cluster pair distribution function for the lowest temperature and the two densities under consideration. In the inset the corresponding particle-particle intracluster pair distribution functions are shown. Our previous discussion concerning the average intercluster potential (cf. right graph in Figure 1) explains that clusters are separated by the correlation length $\lambda_{0}\approx 20\sigma$. The smaller size of the clusters for low density implies the shift of the first peak of $g_{cl-cl}(r)$ to lower distances. The splitting of the second coordination shell for $\rho\sigma^{3}=0.05$ is a typical signature of the presence of an amorphous structurePan et al. (2011), although in our case there is a residual diffusivity. This feature is found in dense amorphous systems, but here the total number density is very low ($\rho\sigma^{3}=0.05$). However, we must bear in mind that the excluded volume of the clusters is much larger than what would correspond to their actual size, as previously discussed (cf. Figure 2). Recall that the effective cluster diameter (corresponding to the excluded volume) is approximately twice the spatial extent of the clusters. On the other hand, once the particle density is halved the splitting disappears and the system becomes more liquid-like, even if the cluster diffusivity remains relatively low ($D/(\sigma^{2}\zeta/m)^{1/2}=0.006$).
@@
If we now move from real to Fourier space, in Figure 10 we have the cluster-cluster structure factor, $S_{cl-cl}(Q)$. A first feature is to be noted, the height of the first peak for both densities is beyond Hansen-Verlet’sHansen and Verlet (1969) freezing rule $S(Q_{max})\gtrsim 2.86$ for which a transition towards a crystal is to be expected. As discussed above, our system presents features of an amorphous solid, but with a somewhat peculiar structure factor. We see that $S(Q)\approx 0$, $\forall\;Q\lesssim 0.2\sigma^{-1}$. In fact, one finds that the hyperuniformity index $H=S(Q_{max})/S(0<Q\lesssim 0.2\sigma^{-1})\approx 10^{3}$ is considerably high. While this value is still away from the criterion for near hyperuniformity, $H\gtrsim 10^{4}$ proposed by Atkinson and coworkers Atkinson et al. (2016) in maximally random jammed systems, it represents a considerable attenuation of density fluctuations. Actually it meets the criterion of effective hyperuniformity, $H\gtrsim 10^{2}$ proposed by Chen and et al.Chen et al. (2018). As to the low-$Q$ limit, it is illustrated in the inset of Figure 10. We appreciate the a linear dependence with $Q^{2}$. Our system, being “effectively hyperuniform”, does not reach the limiting behavior $\lim_{Q\rightarrow 0}S_{cl-cl}(Q)\approx 0$, but we one gets closer as density increases (but this is hardly affected by temperature). This effect is most likely due to the increase of the effective repulsion, since the internal cluster density and average cluster size rise when the overall density is increased. The $Q^{2}$ dependence found in Figure 10 is consistent with a class I (effectively) hyperuniform material Torquato (2018). Local number fluctuations will further confirm this result as shown below.
Now a few words concerning the position of the first maximum of the structure factor, $Q_{max}$. We see that for $\rho\sigma^{3}=0.05$, $Q_{max}\sim 2\pi/\lambda_{0}$ whereas $Q_{max}(\rho\sigma^{3}=0.025)>2\pi/\lambda_{0}$. This reflects the smaller size of the clusters for the lower density (see the corresponding radii of gyration in Figure 5) and matches the shift to lower $r$ of the first maximum of $g_{cl-cl}$ in Figure 9. The quantity $\lambda_{0}$ is sort of a “ground state” estimate of the cluster size, as temperature rises or density decreases its accuracy is lower.
@@
If we now look at the total structure factor, $S(Q)$ depicted in Figure 11, its most noticeable feature is the huge magnitude of the prepeak, which is the signature of a large degree of clustering. Obviously the prepeak positions correspond to that of the first maxima in the cluster-cluster structure factors discussed above. As to the issue of particle hyperuniformity, the situation is less clear. Even if the hyperuniformity index meets the criterion of effective hyperuniformity as defined above, in the inset one can appreciate that the small-Q values of S(Q) are far from small. The results for $\rho\sigma^{3}=0.05$ seem to approach zero as $Q\longrightarrow 0$, but our sample size does not allow for probing smaller $Q$-vectors. The scattering from small clusters and free particles seems to destroy hyperuniformity in a region where otherwise the cluster structure can be deemed as effectively hyperuniform (compare the region for $Q\sigma\lesssim 0.2$ in Figures 11 and 10).
@@
Now, in the inset we can see the large wavenumber behavior of S(Q). The peak maxima in this region $Q_{max}\sigma\approx 8.1$ reflects interparticle correlations at $0.8\sigma$, which is precisely the effective particle diameter, $\sigma_{eff}$. This is in accordance with the maxima of $g(r)$ in the inset of Figure 9. Interestingly, $H=S(Q_{max})/S(Q_{min})=0.2\times 10^{2}$ which is not very far from the criterion of effective hyperuniformity. Note that both $Q_{max}$ and $Q_{min}$ correspond to the intracluster structure factor. A proper choice of the short range interparticle interactions might actually drive the system closer to exhibiting an additional wavenumber range where a considerable attenuation of radiation probes could be present.
As an additional assessment of the effective hyperuniform character of our systems, we have also analyzed the local density (or particle number) fluctuations, defined as
$\sigma^{2}_{N}(R)=<N^{2}>_{R}-<N>_{R}^{2}$ (9)
where $N$ refers to the number of particles and the subscript $R$ refers to the radius of a sampling spherical volume. This quantity is known to fulfillTorquato and Stillinger (2003) $\sigma^{2}_{N}(R)=<N>_{R}$ for a Poisson point pattern, which epitomizes a completely random disordered and uncorrelated system. In fact, in many disordered cases $\sigma^{2}_{N}(R)\propto R^{d}$, where $d$ is the dimensionality ($<N>_{R}\propto R^{d}$ in a uniform system). At the other end of the spectrum, one has point patterns known as hyperuniform that fulfill $\sigma^{2}_{N}(R)\propto R^{d-\alpha}$ with $\alpha>0$. Among these one finds regular patterns such as those of crystals and quasicrystals for which $\alpha=1$Torquato and Stillinger (2003). Other intermediate cases exist (e.g. with logarithmic scaling), with the term strong hyperuniformity reserved for those systems in which $\alpha\geq 1$ Oğuz et al. (2017). Back to our problem, in Figure 12 we present the corresponding analysis based on the configurations of the clusters’ center of mass. The number fluctuation is computed by averaging both the number of clusters, $N_{cl}$ and $N_{cl}^{2}$ for sample spheres of radius $R$ randomly placed in the simulation box. Note that particle number fluctuations are strongly affected by the periodic nature of our sample, since sampling over r-space is limited in accuracy by the Fourier space constraint that must be satisfied in the presence of periodic boundary conditions, namely, ${\bf Q}=(n_{x},n_{y},n_{z})2\pi/L$, where $n_{i}$ are integer numbers and $L$ is the side of the simulation box. For this reason, local number fluctuation can only be accurately computed for values well below half the simulation box size, otherwise effects of the periodic nature of the sample become visible. These limitations have been analyzed in detail by Wright in Ref. Wright (2017). Now, in the left graph of Figure 12, one observes that $\sigma^{2}_{N_{cl}}(R)\propto R^{2}$ except for the high temperature/low density case which exhibits a $R^{3}$ dependence. The curves exhibit strong oscillations, evenly distributed around the linear regression lines (dotted lines in the figure), and indicate that the system has strong spatial correlations consistent with a dense amorphous system. On the right graph the scaled ratio $\sigma^{2}_{N}(R)/(4\pi n_{cl}R^{3}/3)$ is displayed vs the radius of the sampling sphere, $R$. Here $n_{cl}=<N_{cl}>/L^{3}$ is the average density of clusters. A decreasing trend in the $R$-dependence of this ratio is a signature of hyperuniformity Torquato (2018). Both graphs in Figure 12 indicate that our systems (with the exception of the low density/high T case) display inter-cluster structures that can be deemed effective hyperuniform from the point of view of number (or local density) fluctuations. When the right graph of Figure 12 is compared with Figure 2 of Ref. Torquato (2018), we can clearly appreciate that our systems are somewhere in between ordered and disordered hyperuniform point configurations, as could be expected from a glassy-like system.
Figure 12: Left) Local particle number variation of cluster center of mass configurations vs the sample sphere radius squared, $R^{2}$. The linear dependence is an indication of a hyperuniform configuration. Deviation in linearity is apparent in the high T/low density case. Dotted lines represent linear regressions to the simulation data. Right) Local particle number variation of cluster center of mass configurations scaled with the corresponding random uniform value , $34\pi n_{cl}R^{3}/$ 3 vs the sample radius, $R$. A decaying trend indicates the presence of a disordered hyperuniform system. Again, for the high T/low density case the scaled number fluctuation tends to remain constant.
The same analysis can be performed on particle number fluctuations. These results are presented in similar graphs in Figure S3 of the Supplementary Information. Here the situation is somewhat inconclusive. On one hand, the decreasing trend of $\sigma^{2}_{N}(R)/(4\pi\rho R^{3}/3)$ vs $R$ might be an indication of the presence of a certain degree of hyperuniformity. On the other, the strong oscillations on the fluctuation itself (left graph of S3) prevent to reach a clear conclusion. Most likely, the traces of hyperuniformity result from the underlying effective hyperuniform structure of the clusters’ centers of mass.
One might conclude that from the point of view of cluster configurations, the systems studied (with the exception of the high temperature/low density one) can be cast into class I of effectively hyperuniform materials, since $\sigma^{2}_{N}(R)\sim R^{d-1}$. This corresponds to $\lim_{Q\rightarrow 0}S(Q)\propto Q^{\alpha}$ Torquato (2018) with $\alpha>1$. In our case the low-Q behavior matches $\alpha\sim 2$, even if the structure factors do not vanish completely in the zero wavenumber limit. As to the structure displayed by the particle fluid as a whole, no clear conclusion can be drawn at this stage.
@@ VI Conclusions
In summary, we have presented a study of the structural and dynamic properties of a simple model of a type-III SALR self-associating fluid, focusing on the globular cluster phase that upon aggregation displays a cluster structure with effective hyperuniform disorder. Here we have studied both cluster size distributions, cluster density profiles, diffusion and frequency spectra as well as pair distributions (inter and intracluster) and structure factors. From the latter, we have found that the cluster phase for moderate/high cluster densities ($\rho_{C}\gtrsim 0.8$) approaches a cluster glassy state. In these conditions the cluster structure shows a considerable attenuation of density fluctuations for a range of wavenumbers below $0.2\sigma^{-1}$, meeting the criteria of effective hyperuniformity and stealthiness. When considering the system as a whole, which includes contributions from free particles, and effects of cluster size polydispersity, results are less conclusive. Probably larger samples might throw some light onto this problem, but they are beyond our current computational capabilities. One must bear in mind anyway that a complete study must incorporate the explicit influence of particle form factors, which play a fundamental role in determining the interaction with radiation probes. Using as starting point our glassy cluster fluid, for large colloidal particles ($\sigma\approx 1\sim 10\mu m$), attenuation would take place for radiation in the low frequency range of radio waves. The region of interest can be tuned resorting to mixtures of cluster-forming particles Chen et al. (2018) and/or controlling the size of the clusters. Future work will focus on mixtures where intracluster phase separation can enable the tuning of anisotropic cluster-cluster interactions and systems in which self-limitation is controlled by saturation of associative sites in order to reduce polydispersity.
Supplementary material
As supplementary information a graphic illustration of the aggregation process is included. In addition, velocity self-correlation functions are discussed an the net particle number fluctuation as a function of sampling radius is also introduced to complement the analysis in the Section V.
Acknowledgments
The authors acknowledge the support from the Agencia Estatal de Investigación and Fondo Europeo de Desarrollo Regional (FEDER) under grants No. PID2023-151751NB-I00 and PID2020-115722GB-C22. We also would like to acknowledge the Galicia Supercomputing Center (CESGA) for the access to their computer facilities.